### v1

@@ Abstract
As a significant application of multi-source information fusion in intelligent transportation perception systems, Referring Multi-Object Tracking (RMOT) involves localizing and tracking specific objects in video sequences based on language references. However, existing RMOT approaches often treat language descriptions as holistic embeddings and struggle to effectively integrate the rich semantic information contained in language expressions with visual features. This limitation is especially apparent in complex scenes requiring comprehensive understanding of both static object attributes and spatial motion information. In this paper, we propose a Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework that addresses these challenges. It adapts the "what" and "where" pathways from human visual processing system to RMOT tasks. Specifically, our framework comprises three collaborative components: (1)The Bidirectional Interactive Fusion module first establishes cross-modal connections while preserving modality-specific characteristics; (2) Building upon this foundation, the Progressive Semantic-Decoupled Query Learning mechanism hierarchically injects complementary information into object queries, progressively refining object understanding from coarse to fine-grained semantic levels; (3) Finally, the Structural Consensus Constraint enforces bidirectional semantic consistency between visual features and language descriptions, ensuring that tracked objects faithfully reflect the referring expression. Extensive experiments on different benchmark datasets demonstrate that CDRMT achieves substantial improvements over state-of-the-art methods, with average gains of 6.0% in HOTA score on Refer-KITTI and 3.2% on Refer-KITTI-V2. Our approach advances the state-of-the-art in RMOT while simultaneously providing new insights into multi-source information fusion.
@@ 1 Introduction
Therefore, we propose Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework, as illustrated in Figure 1-C. The key innovation of CDRMT lies in its biologically inspired architecture that imitates human visual processing system. Specifically, we first propose a Bidirectional Interactive Fusion (BIF) module to handle cross-modal interactions. As the retinal ganglion cells perform initial sensing and filtering of visual signals before transmission to higher brain regions, BIF (Figure 1-C-Highlight 1) processes visual and language features before the main encoder, allowing each modality to selectively focus on relevant information while preserving its unique characteristics. Moreover, we design a Progressive Semantic-Decoupled Query Learning (PSDQL) module, as depicted in Figure 1-C-Highlight 2. Similar to the human visual system processes information through ventral and dorsal streams, PSDQL progressively injects decoupled semantic information into queries via two complementary pathways. This hierarchical injection enables queries to better capture both static and dynamic characteristics of referred objects. Finally, we introduce the Structural Consensus Constraint (SCC) mechanism, which serves as an additional supervisory signal during training to ensure comprehensive understanding of object appearance and language descriptions. Analogous to the higher cortical areas in the human brain that send feedback signals to earlier processing regions to refine perception and ensure consistency between prediction and sensory input. SCC (Figure 1-C-Highlight 3) employs a text decoder to reconstruct original descriptions from encoded visual features and computes reconstruction loss between reconstructed and original expressions. Extensive experiments on multiple benchmark datasets demonstrate that CDRMT achieves substantial performance improvements, with average gains of 6.0% in HOTA score on Refer-KITTI and 3.2% on Refer-KITTI-V2. The main contributions of this work can be summarized as follows:
@@
  • We introduce a Structural Consensus Constraint that enforces bidirectional semantic consistency between reconstructed and original expressions.
The rest of this paper is organized as follows. Section 2 provides a comprehensive review of related work in visual object tracking. Section 3 details our CDRMT methodology and architectural components. Section 4 presents experimental evaluations and comparative analyses. Section 5 examines current limitations and explores future research directions, with the conclusion provided in Section 6.
@@ 2.1.1 Single-Object Tracking
Visual object tracking has witnessed significant progress over the past decades. Early success was achieved by Siamese networks [4] and discriminative correlation filters [44, 45, 63], which established fundamental tracking paradigms. SiamRPN [31] revolutionized the field by introducing region proposal network (RPN) from object detection into the Siamese tracking framework, while its successor SiamRPN++ [30] further enhanced performance by incorporating an effective sampling strategy to address spatial invariance limitations. However, these template-matching methods, which rely on fixed template frames, struggle with appearance variations caused by occlusion and deformation. Online trackers [5, 69, 75] address this limitation by adaptively updating object templates under specific conditions, leading to improved robustness. The advent of transformers[42, 17] brought another paradigm shift, with TransT [11] pioneering the use of transformer backbones in tracking tasks. Subsequently, OSTrack [65] achieved superior performance by leveraging MAE-pretrained ViT and candidate elimination strategies. GRM [21] further advanced the field by introducing adaptive token partition techniques, effectively addressing the object-background confusion problem prevalent in existing tracking algorithms. Despite these advances in single-object tracking, real-world scenarios often involve multiple objects of interest, highlighting the limitations of SOT methods in handling complex scenes with multiple interactive objects.
@@ 2.1.2 Multi-Object Tracking
Multi-Object Tracking (MOT), which aims to simultaneously detect and track multiple moving objects while maintaining consistent identities across frames. MOT has become crucial for applications ranging from video surveillance to autonomous driving and sports analysis. Two primary paradigms have emerged in MOT: Tracking-by-Detection [58] and Joint-Detection-and-Tracking [71]. The former separates detection and association into two stages, while the latter integrates them in an end-to-end manner. Various specialized benchmarks have driven progress in different scenarios, from pedestrian tracking [20] (PETS2009) to more challenging domains like autonomous driving [22, 67] (KITTI, BDD100K), sports analysis [16] (SportsMOT), and aerial tracking [6] (VisDrone). Recent transformer-based approaches [79, 68, 73] have achieved state-of-the-art performance by better modeling long-range dependencies and complex object interactions. However, conventional MOT methods, which typically track all visible objects in the scene, may not align with practical applications where users are only interested in specific objects.
@@ 2.2.2 Referring Multi-Object Tracking
To address the problem of RSOT, Wu et al. [59] first established a Transformer-based architecture (TransRMOT) that addressed cross-modal alignment and temporal consistency. However, existing approaches often faced scalability limitations when handling complex language queries. Nguyen et al. [50] bridged this gap through their MENDER framework, introducing a Type-to-Track paradigm that employed third-order tensor decomposition to efficiently map language types to tracklet embeddings. He [28] proposed a hierarchical fusion architecture featuring an enhanced early-fusion module, bidirectional cross-modal encoder, and dedicated decoder. Furthermore, Du et al. [18] advanced adaptive feature extraction via its insertable Knowledge Unification Network (iKUN), where textual guidance directly modulates visual feature selection through learnable projection matrices. Recent efforts have also expanded RMOT’s applicability to complex real-world scenarios. Wu et al. [60] contributed the NuPrompt benchmark for autonomous driving, establishing PromptTrack as an end-to-end baseline with 3D-aware language prompts. However, these methods lack effective mechanisms to disentangle and process the complex semantic information contained in referring expressions, which is crucial for accurate referring multi-object tracking. In contrast, our Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework disentangles the rich semantic information in referring expressions, significantly improving tracking performance in complex scenarios.
@@ 3.1 Overview
In this paper, we present the Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework, as shown in Fig 2. Specifically, given the $t$-th frame $\mathbf{F}_{t}$ of video sequence $\mathcal{V}$ and language descriptions $\mathcal{T}$, we extract visual features $\mathbf{V}_{t}\in\mathbb{R}^{H\times W\times C}$ and text features $\mathbf{S}\in\mathbb{R}^{L\times D}$ through pre-trained visual encoder $E_{v}$ and language encoder $E_{l}$, respectively. Here, $H$ and $W$ represent the spatial dimensions of visual features, $L$ is the length of words in semantics, $C$ and $D$ denote the feature channel dimension. Inspired by the human visual processing mechanism, we decompose the text features into two complementary components: static object attributes $\hat{\mathbf{f}}_{so}$ corresponding to the "what" pathway, and spatial motion attributes $\hat{\mathbf{f}}_{sm}$ corresponding to the "where" pathway. We propose a progressive semantic-decoupled query learning (Fig.2-A) mechanism that the static object attributes $\hat{\mathbf{f}}_{so}$ are first injected into queries $\mathbf{Q}_{\text{so}}$ to localize candidate objects based on appearance features. Subsequently, the spatial motion attributes $\hat{\mathbf{f}}_{sm}$ are incorporated into $\mathbf{Q}_{\text{sm}}$ to refine the query candidates, better aligning them with the referred objects in the descriptions. Finally, we introduce a text decoder $\mathcal{D}_{t}$ that reconstructs the original descriptions from visual features $\mathbf{F}^{v}_{t}$, and employ the Structural Consensus Constraint (SCC) mechanism to compute reconstruction loss between the generated and original expressions, as illustrated in Fig.2-B. It serves as an additional constraint that enforces bidirectional semantic consistency between visual objects and language descriptions, ensuring the model’s comprehensive understanding of both object appearance and linguistic attributes.
Figure 2: The architecture of our proposed Cognitive Disentanglement for Referring Multi-Object Tracking framework. It processes sequential video frames through visual and text encoders, followed by bidirectional interactive fusion (BIF) and hierarchical query semantic injection to enhance referring multi-object tracking. (A) The Progressive Semantic-Decoupled Query Learning (PSDQL) module, inspired by the dual-stream ("what"/"where" ) processing mechanism in human visual system, separates language into static object attributes ("what" pathway) and spatial motion information ("where" pathway) to guide object queries. (B) The Structural Consensus Constraint (SCC) mechanism is only applied during the training stage, which enforces geometric consistency between original text embeddings and their reconstructed counterparts to enhance semantic alignment between visual objects and linguistic descriptions.
@@ 3.2 Multimodal Feature Encoding
Figure 3: Overview of the Bidirectional Interactive Fusion.
@@
$\tilde{\mathbf{V}}^{\prime}_{v}$ $=\tilde{\mathbf{V}}_{t}+\text{LN}(\text{CA}(\tilde{\mathbf{V}}_{t},\tilde{\mathbf{S}},\tilde{\mathbf{S}})),$
@@
where ${W}_{q},{W}_{k},{W}_{v}$ are learnable projection matrices. The cross-attended features are further refined through feed-forward networks with residual connections:
$\hat{\mathbf{V}}_{t}$ $=\tilde{\mathbf{V}}^{\prime}_{v}+\text{LN}(\text{MLP}(\tilde{\mathbf{V}}^{\prime}_{v}))\in\mathbb{R}^{HW\times d},\ \hat{\mathbf{S}}=\tilde{\mathbf{S}}^{\prime}+\text{LN}(\text{MLP}(\tilde{\mathbf{S}}^{\prime}))\in\mathbb{R}^{L\times d}.$ (5)
This hierarchical structure enables progressive refinement of cross-modal representations while maintaining modality-specific context, facilitating precise alignment between spatial visual features and semantic linguistic elements. Following the integration of visual and linguistic modalities, we employ a stack of deformable encoder layers [79, 59, 74], to further enhance the language-guided visual features:
@@ 3.3 Progressive Semantic-Decoupled Query Learning
Most existing RMOT approaches [59, 74, 28, 47, 34] directly inject language prompts as holistic features into visual features or queries, which overlook the rich semantic information embedded in language descriptions. This plain fusion strategy particularly struggles with complex motion descriptions (e.g., "black cars in the right"), where models often fail to accurately comprehend and localize objects. Inspired by the dual-stream hypothesis in human visual cognition[1, 14, 3], we propose to decouple object perception and context awareness for better language-guided multi-object tracking. Specifically, we decompose referring expressions into static object descriptions (e.g., object category, appearance attributes) and spatial motion descriptions (e.g., motion states, spatial relationships). This decoupling not only aligns with human visual cognitive mechanisms but also enables hierarchical object localization. We propose a semantic decomposition mechanism to effectively parse the input language expression $\mathcal{T}$ into two complementary streams:
@@
For instance, given the expression “black cars in the right”, the parser would decompose it into static cues $\mathcal{T}_{so}$: “black cars” focusing on static object attributes, and spatial motion information $\mathcal{T}_{sm}$: “in the right” describing spatial relationships. Moreover, to maintain the holistic semantic understanding, we augment both feature streams by incorporating the global sentence embedding $\mathbf{S}_{g}$, which provides essential contextual grounding for the decomposed components. It enables our model to process different semantic aspects more effectively while preserving their inherent relationships within the original expression. Formally:
@@
$\hat{\mathbf{f}}_{so}$ $=\text{LN}([\mathbf{f}_{so}+\mathbf{S}_{g};\mathbf{f}_{so}])\in\mathbb{R}^{K_{o}\times d},$ (9)
$\hat{\mathbf{f}}_{sm}$ $=\text{LN}([\mathbf{f}_{sm}+\mathbf{S}_{g};\mathbf{f}_{sm}]\in\mathbb{R}^{K_{m}\times d},$
where $\mathcal{F}_{\text{lan}}(\cdot)$ denotes a pre-trained language model [41], $K_{o}$ and $K_{m}$ represent the lengths of different attributes in semantics respectively, $D$ and $d$ are the feature dimension, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] indicates feature concatenation in channel dimension, and $\text{LN}(\cdot)$ applies layer normalization to stabilize feature distributions.
Figure 4: Overview of the Query Semantic Injection (QSI) module architecture. It efficiently incorporates disentangled semantic features ($f_{so}$ for static object attributes or $f_{sm}$ for spatial motion information) into query representations $Q_{t}$. The Attention Active Module (AAM) dynamically enhances feature saliency through multi-perspective feature aggregation, generating adaptive attention weights that modulate the cross-attention process. This enables queries to selectively assimilate significant information, facilitating fine-grained semantic-aware object localization.
@@
Specifically, our approach integrated semantic information into detection and tracking query sets $\mathbf{Q}_{t}=\{{\mathbf{Q}^{Det}_{t},\mathbf{Q}^{Tra}_{t}}\}$ through progressive learning. It creates a synergistic relationship between linguistic cues and visual representations, circumventing potential semantic dilution during feature propagation and enabling fine-grained object localization through semantic pattern matching. Our decoder architecture comprises $L$ transformer layers structured hierarchically. The initial $L/2$ layers establish foundational object perception capabilities, while the subsequent layers develop sophisticated contextual understanding. To formalize this process, we introduce the Query Semantic Injection (QSI) Module, which incorporates semantic features into query representations during the transformer decoding phase, as shown in Figure 4. To maximize the utilization of important information in each vector, we employ the Attention Active Module (AAM) to enhance the object information in the attention vector. QSI first enhances queries $\mathbf{Q}_{t}$ by integrating static semantic features $\hat{\mathbf{f}}_{so}$ through the following formulation:
$AAM(\textbf{X})$ $=\textbf{X}+\textbf{X}\odot\mathcal{W}(\textbf{X}),$ (10)
@@
where, $\hat{\mathbf{f}}_{so}$ denotes fundamental object characteristics, including appearance, category, and chromatic properties. $\Psi_{1}$ and $\Psi_{2}$ represent feature transformation and fusion MLPs respectively, $\sigma$ denotes the Sigmoid activation function, $\odot$ indicates element-wise multiplication and $\text{LN}(\cdot)$ is the layer normalization function. The resulting queries $\hat{\mathbf{Q}}^{so}_{t}$ are then fed into decoder layers ($\text{DecLayers}(\cdot)$) for interaction with visual features:
$\mathbf{Q}^{\prime}_{t}$ $=\text{DecLayers}_{0{\rightarrow}L/2}(\mathbf{Q}^{so}_{t},\mathbf{F}^{v}_{t}).$ (12)
@@
3.4 Structural Consensus Constraint
Referring Multi-Object Tracking (RMOT) aims to simultaneously track and associate multiple objects in videos based on natural language descriptions. While existing RMOT methods [59, 74, 9] have shown promising results by leveraging cross-modal features, they often struggle in challenging scenarios with multiple similar objects or complex linguistic descriptions. This misalignment stems from a fundamental challenge: the mapping between visual objects and language descriptions is inherently many-to-many—the same object can be described differently from various perspectives (e.g., "left cars in light color" or "left cars in silver"), while similar descriptions might refer to different objects. Previous works typically do not fully capture these complex semantic relationships and lack effective mechanisms to verify whether tracked objects truly correspond semantically to the referring expressions.

### v3

@@ Abstract
As a significant application of multi-source information fusion in intelligent transportation perception systems, Referring Multi-Object Tracking (RMOT) involves localizing and tracking specific objects in video sequences based on language references. However, existing RMOT approaches often treat language descriptions as holistic embeddings and struggle to effectively integrate the rich semantic information contained in language expressions with visual features. This limitation is especially apparent in complex scenes requiring comprehensive understanding of both static object attributes and spatial motion information. In this paper, we propose a Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework that addresses these challenges. It adapts the "what" and "where" pathways from the human visual processing system to RMOT tasks. Specifically, our framework first establishes cross-modal connections while preserving modality-specific characteristics. It then disentangles language descriptions and hierarchically injects them into object queries, refining object understanding from coarse to fine-grained semantic levels. Finally, we reconstruct language representations based on visual features, ensuring that tracked objects faithfully reflect the referring expression. Extensive experiments on different benchmark datasets demonstrate that CDRMT achieves substantial improvements over state-of-the-art methods, with average gains of 6.0% in HOTA score on Refer-KITTI and 3.2% on Refer-KITTI-V2. Our approach advances the state-of-the-art in RMOT while simultaneously providing new insights into multi-source information fusion.
@@ 1 Introduction
Therefore, we propose Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework, as illustrated in Figure 1-C. The key innovation of CDRMT lies in its biologically inspired architecture that imitates human visual processing system. Specifically, we first propose a Bidirectional Interactive Fusion (BIF) module to handle cross-modal interactions. As the retinal ganglion cells perform initial sensing and filtering of visual signals before transmission to higher brain regions, BIF (Figure 1-C-Highlight 1) processes visual and language features before the main encoder, allowing each modality to selectively focus on relevant information while preserving its unique characteristics. Moreover, we design a Progressive Semantic-Decoupled Query Learning (PSDQL) module, as depicted in Figure 1-C-Highlight 2. Similar to the human visual system processes information through ventral and dorsal streams, PSDQL progressively injects decoupled semantic information into queries via two complementary pathways. This hierarchical injection enables queries to better capture both static and dynamic characteristics of referred objects. Finally, we introduce the Structural Consistency Constraint (SCC) mechanism, which serves as an additional supervisory signal during training to ensure comprehensive understanding of object appearance and language descriptions. Analogous to the higher cortical areas in the human brain that send feedback signals to earlier processing regions to refine perception and ensure consistency between prediction and sensory input. SCC (Figure 1-C-Highlight 3) employs a text decoder to reconstruct original descriptions from encoded visual features and computes reconstruction loss between reconstructed and original expressions. Extensive experiments on multiple benchmark datasets demonstrate that CDRMT achieves substantial performance improvements, with average gains of 6.0% in HOTA score on Refer-KITTI and 3.2% on Refer-KITTI-V2. The main contributions of this work can be summarized as follows:
@@
  • We introduce a Structural Consistency Constraint that enforces bidirectional semantic consistency between reconstructed and original expressions.
The rest of this paper is organized as follows. Section 2 provides a comprehensive review of related work in visual object tracking. Section 3 details our CDRMT methodology and architectural components. Section 4 presents experimental evaluations and comparative analyses. Section 5 offers visualization analyses. Section 6 examines current limitations and explores future research directions, with the conclusion provided in Section 7.
@@ 2.1.1 Single-Object Tracking
Visual object tracking has witnessed significant progress over the past decades. Early success was achieved by Siamese networks [4] and discriminative correlation filters [44, 45, 64], which established fundamental tracking paradigms. FBACF [70] proposed a Feature Block-Aware Correlation Filter for UAV tracking that addresses the limitations of conventional DCF-based methods. SiamRPN [30] revolutionized the field by introducing region proposal network (RPN) from object detection into the Siamese tracking framework, while its successor SiamRPN++ [29] further enhanced performance by incorporating an effective sampling strategy to address spatial invariance limitations. However, these template-matching methods, which rely on fixed template frames, struggle with appearance variations caused by occlusion and deformation. Online trackers [5, 73, 78] address this limitation by adaptively updating object templates under specific conditions, leading to improved robustness. The advent of transformers [42, 15] brought another paradigm shift, with TransT [10] pioneering the use of transformer backbones in tracking tasks. Subsequently, OSTrack [66] achieved superior performance by leveraging MAE-pretrained ViT and candidate elimination strategies. GRM [19] further advanced the field by introducing adaptive token partition techniques, effectively addressing the object-background confusion problem prevalent in existing tracking algorithms. ETDMT [72] developed an Efficient Template Distinction Modeling Tracker with Temporal Contexts that explicitly differentiates between real target and background areas in the template. Despite these advances in single-object tracking, real-world scenarios often involve multiple objects of interest, highlighting the limitations of SOT methods in handling complex scenes with multiple interactive objects.
@@ 2.1.2 Multi-Object Tracking
Multi-Object Tracking (MOT), which aims to simultaneously detect and track multiple moving objects while maintaining consistent identities across frames. MOT has become crucial for applications ranging from video surveillance to autonomous driving and sports analysis. Two primary paradigms have emerged in MOT: Tracking-by-Detection [59] and Joint-Detection-and-Tracking [75]. The former separates detection and association into two stages, while the latter integrates them in an end-to-end manner. Various specialized benchmarks have driven progress in different scenarios, from pedestrian tracking [18] (PETS2009) to more challenging domains like autonomous driving [20, 68] (KITTI, BDD100K), sports analysis [14] (SportsMOT), and aerial tracking [6] (VisDrone). Recent transformer-based approaches [83, 69, 76, 71] have achieved state-of-the-art performance by better modeling long-range dependencies and complex object interactions. However, conventional MOT methods, which typically track all visible objects in the scene, may not align with practical applications where users are only interested in specific objects.
@@ 2.2.2 Referring Multi-Object Tracking
To address the problem of RSOT, Wu et al. [60] first established a Transformer-based architecture (TransRMOT) that addressed cross-modal alignment and temporal consistency. However, existing approaches often faced scalability limitations when handling complex language queries. Nguyen et al. [49] bridged this gap through their MENDER framework, introducing a Type-to-Track paradigm that employed third-order tensor decomposition to efficiently map language types to tracklet embeddings. He [27] proposed a hierarchical fusion architecture featuring an enhanced early-fusion module, bidirectional cross-modal encoder, and dedicated decoder. Furthermore, Du et al. [16] advanced adaptive feature extraction via its insertable Knowledge Unification Network (iKUN), where textual guidance directly modulates visual feature selection through learnable projection matrices. Recent efforts have also expanded RMOT’s applicability to complex real-world scenarios. Wu et al. [61] contributed the NuPrompt benchmark for autonomous driving, establishing PromptTrack as an end-to-end baseline with 3D-aware language prompts. Referring Video Object Segmentation (RVOS) [80, 50, 32] has also significantly contributed to RMOT development. Wang et al. [36] introduced Local-Global Context Aware Transformer for language-guided video segmentation, enhancing cross-modal interaction with multi-scale contextual information. However, these methods lack effective mechanisms to disentangle and process the complex semantic information contained in referring expressions, which is crucial for accurate referring multi-object tracking. In contrast, our Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework disentangles the rich semantic information in referring expressions, significantly improving tracking performance in complex scenarios.
@@ 3.1 Overview
In this paper, we present the Cognitive Disentanglement for Referring Multi-Object Tracking (CDRMT) framework, as shown in Figure 2. Specifically, given the $t$-th frame $\mathbf{F}_{t}$ of video sequence $\mathcal{V}$ and language descriptions $\mathcal{T}$, we extract visual features $\mathbf{V}_{t}\in\mathbb{R}^{H\times W\times C}$ and text features $\mathbf{S}\in\mathbb{R}^{L\times D}$ through pre-trained visual encoder $E_{v}$ and language encoder $E_{l}$, respectively. To effectively align language descriptions with visual features while preserving their inherent characteristics, we propose a Bidirectional Interactive Fusion (Figure 2-A) module that enables structured cross-modal interactions before the main encoder stage. Inspired by the human visual processing mechanism, we decompose the text features into two complementary components: static object attributes $\hat{\mathbf{f}}_{so}$ corresponding to the "what" pathway, and spatial motion attributes $\hat{\mathbf{f}}_{sm}$ corresponding to the "where" pathway. We propose a progressive semantic-decoupled query learning (Figure 2-B) mechanism that the static object attributes $\hat{\mathbf{f}}_{so}$ are first injected into queries $\mathbf{Q}_{\text{so}}$ to localize candidate objects based on appearance features. Subsequently, the spatial motion attributes $\hat{\mathbf{f}}_{sm}$ are incorporated into $\mathbf{Q}_{\text{sm}}$ to refine the query candidates, better aligning them with the referred objects in the descriptions. Finally, we introduce a text decoder $\mathcal{D}_{t}$ that reconstructs the original descriptions from visual features $\mathbf{F}^{v}_{t}$, and employ the Structural Consistency Constraint (SCC) mechanism to compute reconstruction loss between the generated and original expressions, as illustrated in Figure 2-C. It serves as an additional constraint that enforces bidirectional semantic consistency between visual objects and language descriptions, ensuring the model’s comprehensive understanding of both object appearance and linguistic attributes.
Figure 2: The architecture of our proposed Cognitive Disentanglement for Referring Multi-Object Tracking framework. It processes sequential video frames through three collaborative components: (A)The Bidirectional Interactive Fusion module first establishes cross-modal connections while preserving modality-specific characteristics. (B) The Progressive Semantic-Decoupled Query Learning (PSDQL) module, inspired by the dual-stream ("what"/"where" ) processing mechanism in human visual system, separates language into static object attributes ("what" pathway) and spatial motion information ("where" pathway) to guide object queries. (C) The Structural Consistency Constraint (SCC) mechanism is only applied during the training stage, which enforces geometric consistency between original text embeddings and their reconstructed counterparts to enhance semantic alignment between visual objects and linguistic descriptions.
@@ 3.2 Multimodal Feature Encoding
Figure 3: Overview of the Bidirectional Interactive Fusion module. The"Bidirectional" specifically refers to the cross-modal interaction pattern between visual and language modalities, rather than a forward-backward process.
@@
$\tilde{\mathbf{V}}^{\prime}_{t}$ $=\tilde{\mathbf{V}}_{t}+\text{LN}(\text{CA}(\tilde{\mathbf{V}}_{t},\tilde{\mathbf{S}},\tilde{\mathbf{S}})),$
@@
where ${W}_{q},{W}_{k},{W}_{v}$ are learnable projection matrices, and CA()($\cdot$)( ⋅ ) denotes the cross-attention operation.
The cross-attended features are further refined through feed-forward networks with residual connections:
$\hat{\mathbf{V}}_{t}$ $=\tilde{\mathbf{V}}^{\prime}_{t}+\text{LN}(\text{MLP}(\tilde{\mathbf{V}}^{\prime}_{t}))\in\mathbb{R}^{HW\times d},\ \hat{\mathbf{S}}=\tilde{\mathbf{S}}^{\prime}+\text{LN}(\text{MLP}(\tilde{\mathbf{S}}^{\prime}))\in\mathbb{R}^{L\times d},$ (5)
where MLP is Multi-Layer Perceptron, and d denotes the model modality-specific dimensions.
This hierarchical structure enables progressive refinement of cross-modal representations while maintaining modality-specific context, facilitating precise alignment between spatial visual features and semantic linguistic elements. Following the integration of visual and linguistic modalities, we employ a deformable encoder $\text{DeformEnc}(\cdot)$  [83, 60, 77] to further enhance the language-guided visual features:
@@ 3.3 Progressive Semantic-Decoupled Query Learning
Most existing RMOT approaches [60, 77, 27, 47, 33] directly inject language prompts as holistic features into visual features or queries, which overlook the rich semantic information embedded in language descriptions. This plain fusion strategy particularly struggles with complex motion descriptions (e.g., "black cars in the right"), where models often fail to accurately comprehend and localize objects. Inspired by the dual-stream hypothesis in human visual cognition [1, 13, 3], we propose to decouple object perception and context awareness for better language-guided multi-object tracking. Specifically, we employ a linguistic parser [26, 53] that analyzes the syntactic structure of the referring expression to separate static object attributes from spatial-motion information. For instance, noun phrases and adjective phrases that describe object appearance (e.g., "black car," "woman") are extracted as static object information, while prepositional phrases, directional terms, and motion verbs (e.g., "on the left," "moving away,") are identified as spatial-motion information. Our decoupling process implements a linguistically-informed mapping function based on part-of-speech (POS) analysis and dependency parsing. For semantically ambiguous constructions, we apply hierarchical syntactic constraints that preserve phrasal integrity while ensuring mutually exclusive assignments between the two information streams. This decoupling not only aligns with human visual cognitive mechanisms but also enables hierarchical object localization. We propose a semantic decomposition mechanism to effectively parse the input language expression $\mathcal{T}$ into two complementary streams:
@@
For instance, given the expression “black cars in the right”, the parser would decompose it into static cues $\mathcal{T}_{so}$: “black cars” focusing on static object attributes, and spatial motion information $\mathcal{T}_{sm}$: “in the right” describing spatial relationships. Moreover, to maintain the holistic semantic understanding, we augment both feature streams by incorporating the enhanced sentence embedding $\hat{\mathbf{S}}$, which provides essential contextual grounding for the decomposed components. It enables our model to process different semantic aspects more effectively while preserving their inherent relationships within the original expression. Formally:
@@
$\hat{\mathbf{f}}_{so}$ $=\text{LN}([\mathbf{f}_{so}+{r}(\hat{\mathbf{S}});\mathbf{f}_{so}])\in\mathbb{R}^{K_{o}\times d},$ (9)
$\hat{\mathbf{f}}_{sm}$ $=\text{LN}([\mathbf{f}_{sm}+{r}(\hat{\mathbf{S}});\mathbf{f}_{sm}]\in\mathbb{R}^{K_{m}\times d},$
where $\mathcal{F}_{\text{lan}}(\cdot)$ denotes a pre-trained language model [41], $K_{o}$ and $K_{m}$ represent the lengths of different attributes in semantics respectively, $D$ and $d$ represent the original text encoder output features dimensions and model modality-specific dimensions, respectively, and $r(\cdot)$ denotes the reshape operation, [;][$\cdot$;$\cdot$][ ⋅ ; ⋅ ] indicates feature concatenation in channel dimension, and $\text{LN}(\cdot)$ applies layer normalization to stabilize feature distributions.
Figure 4: Overview of the Query Semantic Injection (QSI) module architecture. It efficiently incorporates disentangled semantic features ($\hat{\mathbf{f}}_{so}$ for static object attributes or $\hat{\mathbf{f}}_{sm}$ for spatial motion information) into query representations $Q_{t}$. The Attention Active Module (AAM) dynamically enhances feature saliency through multi-perspective feature aggregation, generating adaptive attention weights that modulate the cross-attention process. This enables queries to selectively assimilate significant information, facilitating fine-grained semantic-aware object localization.
@@
Specifically, our approach integrated semantic information into detection and tracking query sets $\mathbf{Q}_{t}=\{{\mathbf{Q}^{Det}_{t},\mathbf{Q}^{Tra}_{t}}\}$ through progressive learning. It creates a synergistic relationship between linguistic cues and visual representations, circumventing potential semantic dilution during feature propagation and enabling fine-grained object localization through semantic pattern matching. Our decoder architecture comprises $L$ transformer layers structured hierarchically. The initial $L/2$ layers establish foundational object perception capabilities, while the subsequent layers develop sophisticated contextual understanding. To formalize this process, we introduce the Query Semantic Injection (QSI) Module, which incorporates semantic features into query representations during the transformer decoding phase, as shown in Figure 4. This enables direct semantic conditioning of object queries, which serve as learnable object prototypes crucial for trajectory modeling in tracking tasks. To maximize the utilization of important information in each vector, we employ the Attention Active Module (AAM), which implements multi-perspective feature aggregation by simultaneously utilizing direct feature transformation alongside complementary statistical pooling operations. It allows the model to dynamically modulate attention weights between query features and semantic embeddings, effectively enhancing feature saliency in a context-aware manner. QSI first enhances queries $\mathbf{Q}_{t}$ by integrating static semantic features $\hat{\mathbf{f}}_{so}$ through the following formulation:
$AAM(\textbf{X})$ $=\textbf{X}+\textbf{X}\mathcal{W}(\textbf{X}),$ (10)
@@
where $\hat{\mathbf{f}}_{so}$ denotes fundamental object characteristics, including appearance, category, and chromatic properties. $\Psi_{1}$ and $\Psi_{2}$ represent feature transformation and fusion MLPs respectively, $\sigma$ denotes the sigmoid function and $\text{LN}(\cdot)$ is the layer normalization function. The resulting queries $\hat{\mathbf{Q}}^{so}_{t}$ are then fed into decoder layers ($\text{DecLayers}(\cdot)$) for interaction with visual features:
$\mathbf{Q}^{\prime}_{t}$ $=\text{DecLayers}_{0{\rightarrow}L/2}(\mathbf{Q}^{so}_{t},\mathbf{F}^{v}_{t}),$ (12)
where $L$ represents the total number of decoder layers in our architecture.
@@
3.4 Structural Consistency Constraint
Referring Multi-Object Tracking (RMOT) aims to localize and track specific objects in video sequences based on language references. While existing RMOT methods [60, 77, 8] have shown promising results by leveraging cross-modal features, they often struggle in challenging scenarios with multiple similar objects or complex linguistic descriptions. This misalignment stems from a fundamental challenge: the mapping between visual objects and language descriptions is inherently many-to-many—the same object can be described differently from various perspectives (e.g., "left cars in light color" or "left cars in silver"), while similar descriptions might refer to different objects. Previous works typically do not fully capture these complex semantic relationships and lack effective mechanisms to verify whether tracked objects truly correspond semantically to the referring expressions.